Kevin Briley began a 114 mile bicycle trip to build up stamina for a triathlete competition. Unfortunately, his bicycle chain broke, so he finished the trip walking. The whole trip took 6 hours. If Kevin walks at a rate of 4 miles per hour and rides at 24 miles per hour, find the amount of time he spent on the bicycle.
Thank you for helping me.

Answer by ankor@dixie-net.com(22740)

evin Briley began a 114 mile bicycle trip to build up stamina for a triathlete competition.
Unfortunately, his bicycle chain broke, so he finished the trip walking. The whole trip took 6 hours. If Kevin walks at a rate of 4 miles per hour and rides at 24 miles per hour, find the amount of time he spent on the bicycle.
:
let t = time spent on the bike
The whole trip took 6 hr, therefore
(6-t) = his time walking
:
write a distance equation; dist = speed * time
24t + 4(6-t) = 114
24t + 24 - 4t = 114
24t - 4t = 114 - 24
20t = 90
t = 90/20
t = 4.5 hrs on the bike
:
:
Check this, he must have spent 1.5 hrs walking
24 * 4.5 = 108 mi
4 * 1.5 = 6 mi
----------------
total dist: 114 mi